lovemusic34552:

-3(a2 + a+1) -4(-a2 - a+1)

Vocaloid:

@lovemusic34552 use the distributive property A(B+C+D) = A*B + A*C + A*D

Vocaloid:

Using this same pattern, -3(a2 + a+1) = ?